OpenStudy (anonymous):

Replace ? with =, >, or < to make the statement true. 16 ÷ 4 + 12 ? 4 + 24 ÷ 2 A. = B. < C. >

OpenStudy (anonymous):

it is equal or =

OpenStudy (anonymous):

16/4+12 = 16 4 + 24 ÷ 2 = 26

OpenStudy (anonymous):

The second half is larger.

OpenStudy (anonymous):

@K-K-K-Kiaa that is completely wrong

OpenStudy (anonymous):

How so? 16/4= 4 Then, 4+12= 16.

OpenStudy (anonymous):

16/4+12 = 4+24/2 4+12 = 4+12 16=16

OpenStudy (anonymous):

4+24= 28, then 28/2= 14. Okay,my mistake

OpenStudy (anonymous):

But it still isn't equal (:

OpenStudy (anonymous):

division comes first in every instance

OpenStudy (anonymous):

@GallifreyDaBunny please disregard kiaa. He/she doesn't now how to properly do math

OpenStudy (anonymous):

;-; Ops. Lol I keep pressing the wrong things. I see what I did.

OpenStudy (anonymous):

Look here. This IS a tutoring site. Don't get snappy.People do make mistakes. @Fransisco_w Calm your tits.

OpenStudy (anonymous):

Im just trying to prevent the wrong answer from being taken. A tutor should be HELPING not adding to the confusion. As for my snappy attitude im sorry. @K-K-K-Kiaa

OpenStudy (anonymous):

That does not mean you can insinuate that I'm "not doing math properly." At least I was able to catch my mistake. @Fransisco_w

OpenStudy (anonymous):

Technically speaking you were not doing math properly and yes you caught your mistake after said you were not doing it properly. As i said before sorry for my attitude, I am just trying to help the person asking the question.

OpenStudy (anonymous):

@K-K-K-Kiaa